and the x axis Find the Area and the integral

f(x)=x^{3}-x^{2}-6x and the x axis

Find the Area and the integral
Answer #1

fcx) = x - x - 6x; y=0 A= d I fox)-gex] ldx a 33-X-6X = 0 X (x+2) CX-3) 20 20; X=-2; a=3 -) 1 x - x - 6xl da 8(.x²+x+6x) dx.